OpenStudy (anonymous):

Andy is twice as old as Kate. In 6 years, their ages will total 60. How old is each now?

OpenStudy (anonymous):

A=2K (A+6)+(K+6)=60 A represents Andy's age. K represents Kate's age. Solve the system of equations, using substitution is probably easiest.

mathslover (mathslover):

first of all : @BabeBri304 \[\huge{\textbf{Welcome to openstudy}}\] Let the age of andy ------->x Let the age of Kate-------->2x after 6 years their ages is : x+6 and 2x+6 and given the total of their ages = 60 \[\huge{x+6+2x+6=60}\] \[\huge{x+2x+12=60}\] \[\huge{x+2x=48}\] \[\huge{3x=48}\] \[\huge{x=\frac{48}{3}=16}\] can u find the age of kate ?

OpenStudy (anonymous):

@mathslover You got that backwards. Andy is twice as old as Kate, not vice versa.

OpenStudy (anonymous):

Other than that, though, good explanation :)

OpenStudy (theviper):

Let the ages of Andy & Kate be x & y respectively. then two eq are formed:- x-2y=0 (x+6)+(y+6)=60

mathslover (mathslover):

oops : sorry no problem here i go with the correct one \[\huge{kate=2andy}\] \[\huge{x=2y}\] hence kate = 2y and andy = y u will get y = 16 and x ? can u find this now

OpenStudy (theviper):

Solve it by elimination or substitution method:)

OpenStudy (anonymous):

thank you

OpenStudy (theviper):

yw:)